OpenStudy (anonymous):

Use a technological tool and Euler's method with ten steps of size 0.1 to estimate y(1) for differential equation dy/dx=y-2. Make certain to include supporting screen captures.

OpenStudy (anonymous):

write a small sweet program.. euler's method is just a linear method in short steps, you keep updating your "y" and timestep. fortran, C and python are the programming choices. if you have access, you may want to try matlab or octave

OpenStudy (anonymous):

i dont know whats matlab or octave...i do have internet

OpenStudy (anonymous):

you know, its easy. you can implement it on excel as columns!!!! as you drag it down the next row should get updated!

OpenStudy (anonymous):

no we need to do it by ourself

OpenStudy (anonymous):

yea.. ok. open up the excel.

OpenStudy (anonymous):

and how to use euler's method here

OpenStudy (anonymous):

done have one dude

OpenStudy (anonymous):

you mean by "hand"?> it says use technological tool.

OpenStudy (anonymous):

you do not have Microsoft Excel?

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

ya i meant euler's method by hand and technological tool by internet

OpenStudy (anonymous):

ok. step 1) initial things iteration n \(h=0.1\) = time step. \(f(x,y) =y-1\) -> equation for update

OpenStudy (anonymous):

what are the initial conditions?

OpenStudy (anonymous):

i really dont know

OpenStudy (anonymous):

that must be provided y(0)=?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

it is 1

OpenStudy (anonymous):

that is the initial condition when \(x_0=0\)

OpenStudy (anonymous):

i know y(0) = 1...where did you get x0 = 0 from

OpenStudy (anonymous):

oh ya it is given nm

OpenStudy (anonymous):

we have \(x0=0,\;y0=1\) iter# 0 \(y_1=y_0+hf(x_0,y_0)=1+(0.1)(1−2)=0.9\) \(x_1=x_0+h=0+0.1=0.1\)

OpenStudy (anonymous):

iter# 1 \(y_2=y_1+hf(x_1,y_1)=0.9+0.1(0.9−2)=0.79\) \(x_2=x_1+h=0.1+0.1=0.2\)

OpenStudy (anonymous):

iter# 2 \(y_3=y_2+hf(x_2,y_2)=0.79+0.1(0.79−2)=0.669\) \(x_3=x_2+h=0.2+0.1=0.3\)

OpenStudy (anonymous):

and you keep going till you get to iter# \(n=10\)

OpenStudy (anonymous):

it is so long lol...isnt there any shortcut

OpenStudy (anonymous):

like I said, some programming tool.

OpenStudy (anonymous):

ok lemme go until n=10...thrn call you back...we will discuss on technological tool then

OpenStudy (anonymous):

whats iter

OpenStudy (anonymous):

iter = short for iteration.. btw, your answer should be close to (not exact) analytical solution \(y(1)=2-e\)

OpenStudy (anonymous):

ohk